OpenStudy (jonathan):

In circle A shown below, the measure of ∠BAD is 148°: If mArc BC is 105°, what is mArc CD? 74° 148° 43° 31°

OpenStudy (savannaadams16):

Is there a picture?

OpenStudy (jonathan):

@savannaadams16

OpenStudy (savannaadams16):

The arc of BCD is the same as the angle BAD. So, since arcBC is inside of arcBCD just subtract it. arcBCD - arcBC

OpenStudy (jonathan):

Whats the measure of BCD?

OpenStudy (savannaadams16):

It would be the same measurement as angle BAD

OpenStudy (jonathan):

so c?

OpenStudy (savannaadams16):

Yes, ArcBAD - ArcBC = 43dgrees 148 -105 =43

OpenStudy (jonathan):

thank you! could you help with another one?

OpenStudy (savannaadams16):

Sure. :)

OpenStudy (jonathan):

In circle A shown below, mArc BC is 61° and mArc EF is 76°: What is m∠FDE? 30.5° 38° 76° 68.5°

OpenStudy (jonathan):

@savannaadams16 you there?

OpenStudy (savannaadams16):

Yes, just trying to find the answer myself.

OpenStudy (jonathan):

I think its 68.5

OpenStudy (savannaadams16):

I believe it would be Angle FDE = 1/2(mBC + mEF)

OpenStudy (savannaadams16):

1/2(61+ 76) = 68.5

OpenStudy (savannaadams16):

Do you see how I got that answer?

OpenStudy (jonathan):

yes, thank you again

OpenStudy (savannaadams16):

You're welcome. :)